Ticket Number: HEA-NOC/20040815-1 Ticket Status: OPEN
Ticket Type: unscheduled Resolver:
Ticket Opened: 20040815 09:59 UTC+1 Problem Start: 20040815 06:56 UTC+1
Ticket Update: 20040815 09:59 UTC+1 Problem End: 20040815 06:57 UTC+1
Between 06:56 and 06:57 there were two quick up / downs to UCC's 100M link
20040815 09:56 SOB Issue seems to have resolved its self. No response from
cork. Will contact again on Monday to check for RFO.
This ticket can be monitored at http://www.hea.net/tickets/20040815-1